Carbex, Inc., produces cutlery sets out of high-quality wood and steel. The company makes a Standard set and a Deluxe set and sells them to retail department stores throughout the country. The Standard set sells for $72, and the Deluxe set sells for $87. The variable expenses associated with each set are given below.

Standard Deluxe
Variable production costs $ 21.00 $ 36.00
Sales commissions (21% of sales price) $ 15.12 $ 18.27

The company’s fixed expenses each month are:

Advertising $ 111,000
Depreciation $ 23,500
Administrative $ 66,000

Mary Parsons, the financial vice president, watches sales commissions carefully and has noted that they have risen steadily over the last year. For this reason, she was shocked to find that even though sales have increased, profits for the current month—May—are down substantially from April. Sales, in sets, for the last two months are given below:

Standard Deluxe Total
April 4,600 2,600 7,200
May 1,600 5,600 7,200

Required:

1-a. Prepare contribution format income statement for April.

1-b. Prepare contribution format income statement for May.

3-a. Compute the break-even point in dollar sales for April.

3-b. Would the break-even point in May be higher or lower than the break-even point in April?
Answer #1

1-a. Income statement for April

Standard

Deluxe

Total(in $)

Sales(A)

=4600*$72

=$331200

=2600*$87

=$226200

557400

Less :Expenses

Variable production Cost

=4600*21

=$96600

=2600*$36

=$93600

190200

Sales commission

=4600*15.12

=$69552

=2600*18.27

=$47502

117054

Advertising

111,000

Depreciation

23,500

Administrative

66,000

Total Expenses(B)

507754

Net Income C=(A-B)

$49646

1-b. Income statement for May

Standard

Deluxe

Total(in $)

Sales(A)

=1600*$72

=$115200

=5600*$87

=$487200

602400

Less :Expenses

Variable production Cost

=1600*21

=$33600

=5600*$36

=$201600

235200

Sales commission

=1600*15.12

=$24192

=5600*18.27

=$102312

126504

Advertising

111,000

Depreciation

23,500

Administrative

66,000

Total Expenses(B)

562204

Net Income C=(A-B)

$40196

3-a. Computation of break-even point in dollar sales for April.

Breakeven points in Units=Fixed cost / Sales price per unit- Variable cost per unit

Standard Product

=           Advertising + Depreciation+ Administrative

        Total sales price per unit of Standard product set-variable cost

=111000+23500+66000 = 1.214

    (72-21-15.12) *4600

Breakeven point in Dollars= Sales price per unit X Break even points in Units  

Standard Product=72X 1.214

                           =87.46

Deluxe Product

Breakeven points in Units=Fixed cost / Sales price per unit- Variable cost per unit

                       111000+23500+66000 =$2.356

                                                (87-36-18.27)*2600

Breakeven point in Dollars= Sales price per unit X Break even points in Units   =87 X 2.356 =204.98

Breakeven Point in April

Standard

Deluxe

Breakeven point in Dollars(per unit)

87.46

204.98

No of units

4600

2600

total

402316

532948

3-b. To analyse whether the breakeven point in May higher than or lower to April, we need to calculate the breakeven point for the month of May for both the set.

Breakeven point for may

Breakeven points in Units=Fixed cost / Sales price per unit- Variable cost per unit

Standard Product

=           Advertising + Depreciation+ Administrative

        Total sales price per unit of Standard product set-variable cost

=111000+23500+66000 = 3.49

    ( 72-21-15.12)*1600

Breakeven point in Dollars= Sales price per unit X Break even points in Units  

Standard Product=72X 3.49

                           =251.46

Deluxe Product

Breakeven points in Units=Fixed cost / Sales price per unit- Variable cost per unit

                       111000+23500+66000 =1.093

                                                (87-36-18.27)*5600

Breakeven point in Dollars= Sales price per unit X Break even points in Units   =87 X 1.093 =95.169

May

Standard

Deluxe

Breakeven point in Dollars(per unit)

251.46

95.169

No of units

1600

5600

total

402336

532946.4

April

May

Standard

402316

402336

Deluxe

532948

532946.4

The breakeven point in case of standard set is higher in May as compared to April due to less no. of units.

The breakeven point in case of deluxe set is lower in May as compared to April due to higher no. of units